I. Fill in the remaining cells of the table. Assume perfectly competition. This firm sells its product for $150 TC 200 AFC ATC MC AVC Profit 2 100 20 240 5 24 660 160
Answer #1

TC = TVC + TFC

TFC = TC - TVC

TVC = TC - TFC

AC = TC/Q

AC = AFC + AVC

AFC = TFC/Q

AVC = TVC/Q

MC = \DeltaTC/\DeltaQ

AFC = 24

AFC = TFC/Q

TFC = AFC\timesQ

= 24\times5

= 120

Q P TR =P\timesQ TC TVC TFC AFC ATC MC AVC Profit = TR - TC
1 150 150 200 80 120 120 200 200 80 - 50
2 150 300 220 100 120 60 110 20 50 80
3 150 450 240 120 120 40 80 20 40 210
4 150 600 360 240 120 30 90 120 60 240
5 150 750 500 380 120 24 100 140 76 250
6 150 900 660 540 120 20 110 160 90 240